How much water must be evaporated from 38 ounces of a 4% salt solution to make a 5% salt solution?
I have to "let x = something"
and then set up an equation and solve it.

Answer by stanbon(75887)

How much water must be evaporated from 38 ounces of a 4% salt solution to make a 5% salt solution?
----------------------
Let "x" be the amount of water that must be evaporated.
---------------
EQUATION:
Salt: 0.04*38 - 0*x = 0.05(38-x)
1.52 = 1.9 -0.05x
-0.05x =-0.38
x = 8 ounces (amount of water that must be evaporated)
==================
Cheers,
Stan H.